1, Madsen Motors's bonds have 10 years remaining to maturity. Interest is paid annually; they have a $1,000 par value; the coupon interest rate is 10%; and the yield to maturity is 11%. What is the bond's current market price? Round your answer to the nearest cent.

2,

YIELD TO MATURITY AND FUTURE PRICE

A bond has a $1,000 par value, 8 years to maturity, and a 7% annual coupon and sells for $980.

  1. What is its yield to maturity (YTM)? Round your answer to two decimal places.

       %
  2. Assume that the yield to maturity remains constant for the next 3 years. What will the price be 3 years from today? Do not round intermediate calculations. Round your answer to the nearest cent.  
Answer #1

Question 1

Bond Valuation: The value of bond is the present value of the expected cashflows from the bond,discounted at Yield to Maturity(YTM).

Prima facie, the bond will trade at discount as YTM>coupon rate

Year Cash flow PVAF/PVF@11% Present Value (Cashflow*PVAF/PVF)
1-10 100 5.8892* 588.92
10 1000 0.3522** 352.19

Current Market Price of Bonds = Cashflow*PVAF/PVF

= 588.92+352.19

= $941.11

*PVAF = (1-(1+r)^-n)/r

**PVF = 1 / (1+r)^n

Note: As per Chegg guidelines, when multiple questions are posted, we are mandated do only the first question. So please post the questions individually.